find eqn. of a line passing through intersectin of lines x+y+1=0 and 2x-3y-1=0 and passing through (1,1) whose slope is 2 units

Grade:

1 Answers

vineet chatterjee
9 years ago
this question is to be done by done by the method of family of lines.
so,by this method-(x+y+1)+lambda(2x-3y-1)=0
now x=--2/5 and y=--3/5 are the solutions of the equation x+y+1 =0 and 2x-3y=1.
so,put the values of x and y in the above equation which i have written in the form of lambda.
thus,you will get the value of lambda and so you can get a general equation which passes through the given two lines.